OpenStudy (anonymous):

The midpoint of a line segment with end points (–10, y1) and (–6, 7) is (–8, 6). What is the value of y1?

OpenStudy (anonymous):

(-10+-6)/2 = -16/2=-8 (y+7)/2=6; can you solve for why?

OpenStudy (anonymous):

that is confusing ! I am not good with these questions.

OpenStudy (anonymous):

(y+7)/2=6 multiply by 2 y+7=12 now?

OpenStudy (anonymous):

5?

OpenStudy (nubeer):

yes it's 5 @Amelia.Watts3412

OpenStudy (anonymous):

correct